# CART: Caltech Aerial RGB-Thermal Dataset in the Wild

### v1

@@ 4.2.2 Performance analysis
All RGB-T models outperform the single-modality baselines, with CMNeXt outperforming the next best RGB-T model by 0.026 mIoU (Tab. 3). Overall, incorporating RGB imagery into the segmentation pipeline greatly improved the distinction between land-based stuff classes which can be hard to distinguish in thermal due to thermal crossover. However, these improvements come at a high computational cost and would not currently be suitable for field robotic applications. Excelling in this benchmark would involve improving segmentation performance of the object classes, possibly by leveraging RGB data in unsupervised domain adaptation training schemes, as well as increasing computational efficiency through prudent architectural design choices.
@@ 4.3.2 Image translation metrics
PSNR $\uparrow$
SSIM $\uparrow$
mIoU $\uparrow$
UNIT [41]
Unpaired GAN
12.85
0.421
0.189
MUNIT [27]
Unpaired GAN
16.08
0.459
0.173
Edge-guided RGB-T [35]
Unpaired GAN
12.20
0.436
0.225
Pix2Pix [28]
Paired GAN
20.06
0.547
0.354
Pix2PixHD [71]
Paired GAN
20.05
0.579
0.379
VQGAN [15]
Paired GAN
18.09
0.540
0.388
Palette [57]
Paired Diffusion
11.51
0.399
0.194
@@ 4.3.3 Performance analysis
Paired GANs outperform unpaired GANs and diffusion methods, achieving higher values across all three metrics (Tab. 4). Qualitative assessment (Fig. 6) shows poor translations from unpaired techniques, with most retaining geometric artifacts unique to RGB (ocean waves and shadows) and ignoring relative temperature characteristics. Paired GANs produce results that appear similar to real thermal images but with inconsistent details upon closer inspection. Likewise, the diffusion method produces accurate translations but is not consistent. Overall, RGB-T translation still requires improvements before being able to provide a reliable source of thermal training data from existing RGB datasets. For domain adaptation purposes, we emphasize that maximizing the thermal mIoU score is more important than photometric consistency.
@@ 4.4 Motion Tracking
N1
N2
N3
N4
N5
N6
C1
C2
C3
C4
D1
D2
VINS-Fusion
SLAM
RGB
2.883
7.530
4.036
2.009
4.045
0.834
1.555
3.277
1.566
1.031
0.700
VINS-Fusion
Thermal
9.454
12.26
9.854
2.926
11.00
5.114
7.296
1.513
5.321
2.025
0.879
VINS-Fusion
VIO
RGB
5.458
9.509
6.518
2.121
6.026
1.277
1.555
3.277
1.377
1.031
0.725
VINS-Fusion
Thermal
13.82
12.86
11.99
2.926
13.98
5.194
7.284
10.60
5.321
2.034
0.879
Open-VINS
VIO
RGB
22.80
36.15
37.15
5.165
3.896
3.573
0.475
0.700
Open-VINS
Thermal
14.43
30.64
16.12
2.258
1.562
5.311
1.073
Trajectory length (m)
1533
1326
1253
153
1104
642
310
65
174
137
88
80
@@ 4.5.1 Zero-shot foundation models on thermal imagery
Current foundation models work well on RGB imagery but have not been tested in the thermal domain, especially in non-urban environments. To evaluate foundation models on thermal imagery, specifically for zero-shot segmentation and semantic segmentation, we compare RGB and thermal performance using the aligned set (Sec. 3.2.4). With this set, we isolate modality as the root cause for any performance difference. In the following experiments, our metrics do not include the rare object classes (vehicle and person) as they drop metrics to near-uniform values for both modalities, making it difficult to compare.
@@
[…]
Thermal
0.018
0.007
0.022
0.204
Color (grayscale)
0.538
0.536
0.531
0.601
Color (col. jitter)
0.800
0.792
0.801
0.841
${}^{\dagger}$refers to AP@0.5::0.95, following MSCOCO [39]
@@
Method Color
Color (gray)
Color (jitter)
Thermal
SAM [33] + GT ${}^{\ddagger}$
0.704
0.698
0.700
0.653
Grounded SAM [42]
0.380
0.351
0.361
0.193
OV-Seg [37]
0.362
0.340
0.365
0.217
ODISE [75]
0.504
0.450
0.486
0.232
${}^{\ddagger}$ Most frequent ground truth class label in a SAM mask assigned as label for entire mask
We quantify Segment Anything’s (SAM) [33] zero-shot segmentation sensitivity to modality change by measuring the spatial alignment of its mask outputs. We apply it to the RGB set and the 8-bit thermal image set (Sec. 3.2.2) with the default 32$\times$32 grid points to segment everything (Fig. 7). We compute the average precision (AP) of the predicted thermal masks, using the predicted RGB masks as ground truth, and compare the thermal AP against the APs of a color-jittered RGB set and a grayscale set. Results show that SAM is sensitive to photometric changes, with sizable performance drops from color-jitter to grayscale to thermal (Tab. 6(a)).
@@
Finally, we create semantic SAM masks by assigning class labels to SAM instances based on the most frequent ground truth class within an instance (Fig. 7). We find that semantic SAM masks attain much higher mIoU score (0.653) compared to the other zero-shot methods. Based on the overall results (Tab. 6), we conclude that SAM can delineate object boundaries in the thermal domain reasonably well but can oversegment instances resulting in poor AP scores when compared to RGB (which can also oversegment). When paired with ground truth semantic labels, this behavior is hidden.
@@ 4.5.2 Transfer learning with urban thermal datasets
mIoU # Params. (M)
None
ImageNet
Therm. Urban
@@
0.640
0.690
0.688
1.1
EfficientViT
0.687
0.725
0.714
4.8
ResNet18
0.702
0.713
0.706
12.3
ResNet50
0.682
0.713
0.728
26.7
ConvNext-B
0.625
0.717
0.697
89.4
@@
[…]
@@ 5 Conclusion
We presented the Aerial RGB-Thermal Dataset, the first publicly available dataset specifically tailored towards advancing thermal semantic perception and motion tracking algorithms in natural environments. We created four distinct benchmarks encompassing semantic segmentation in both thermal and RGB-T domains, RGB-T image translation, and motion tracking, and use these benchmarks to demonstrate the current challenges faced by thermal-based robot perception and localization algorithms. Current semantic segmentation and image translation methods were particularly affected by geographical domain shifts, reflecting the diverse intra-class and inter-scenery distributions in our dataset, as well as temporal domain shifts due to thermal inversion and crossover. Motion tracking algorithms, on the other hand, suffered from poor feature tracking in the thermal modality due to reduced spatial quality as compared to RGB and encountered failures when faced with challenging, feature-sparse scenarios. However, such scenarios are commonly encountered during real-world deployments with failures of significant cost and consequence [1]. Our motion tracking benchmark greatly penalizes current VIO/SLAM algorithms for assuming ideal feature-tracking conditions, serving as a unique testbed for algorithm improvements in this area. This dataset, along with the benchmarks, can be used into the future to help develop algorithms that help expand the operating domains of both robotics and computer vision in general.

### v2

@@ 4.2.2 Performance analysis
All RGB-T models outperform the single-modality baselines, with CMNeXt outperforming the next best RGB-T model by 0.026 mIoU (Tab. 4). Overall, incorporating RGB imagery into the segmentation pipeline greatly improved the distinction between land-based stuff classes which can be hard to distinguish in thermal due to thermal crossover. However, these improvements come at a high computational cost and would not be suitable for field robotic applications. Excelling in this benchmark would involve improving the segmentation performance of the object classes, possibly by leveraging RGB data in unsupervised domain adaptation training schemes.
@@ 4.3.2 Image translation metrics
PSNR $\uparrow$ SSIM $\uparrow$ mIoU $\uparrow$
UNIT [42] Unpaired GAN 12.85 0.421 0.189
MUNIT [27] Unpaired GAN 16.08 0.459 0.173
Edge-guided RGB-T [35] Unpaired GAN 12.20 0.436 0.225
Pix2Pix [28] Paired GAN 20.06 0.547 0.354
Pix2PixHD [73] Paired GAN 20.05 0.579 0.379
VQGAN [15] Paired GAN 18.09 0.540 0.388
Palette [58] Paired Diffusion 11.51 0.399 0.194
@@ 4.3.3 Performance analysis
Paired GANs outperform unpaired GANs and diffusion methods, achieving higher values across all three metrics (Tab. 5). Qualitative assessment (Fig. 6) shows poor translations from unpaired techniques, with most retaining geometric artifacts unique to RGB (ocean waves and shadows) and ignoring relative temperature characteristics. Paired GANs produce results similar to real thermal images but with inconsistent details upon closer inspection. Likewise, the diffusion method produces accurate translations but is not consistent. Overall, RGB-T translation still requires improvements before being able to provide a reliable source of thermal training data from existing RGB datasets. For domain adaptation purposes, we emphasize that maximizing the thermal mIoU score is more important than photometric consistency.
@@ 4.4 Motion Tracking
N1 N2 N3 N4 N5 N6 C1 C2 C3 C4 D1 D2
VINS-Fusion SLAM RGB 2.883 7.530 4.036 2.009 4.045 0.834 1.555 3.277 1.566 1.031 0.700
VINS-Fusion Thermal 9.454 12.26 9.854 2.926 11.00 5.114 7.296 1.513 5.321 2.025 0.879
VINS-Fusion VIO RGB 5.458 9.509 6.518 2.121 6.026 1.277 1.555 3.277 1.377 1.031 0.725
VINS-Fusion Thermal 13.82 12.86 11.99 2.926 13.98 5.194 7.284 10.60 5.321 2.034 0.879
Open-VINS VIO RGB 22.80 36.15 37.15 5.165 3.896 3.573 0.475 0.700
Open-VINS Thermal 14.43 30.64 16.12 2.258 1.562 5.311 1.073
Trajectory length (m) 1533 1326 1253 153 1104 642 310 65 174 137 88 80
@@ 4.5.1 Zero-shot foundation models on thermal imagery
Current foundation models work well on RGB imagery but have not been benchmarked on thermal, especially in non-urban settings. To evaluate these models on thermal imagery, specifically for zero-shot segmentation and semantic segmentation, we compare RGB and thermal performance using the aligned set (Sec. 3.2.4) in order to isolate modality as the root cause for performance differences. In the following tests, our metrics do not include the rare object classes (vehicle, person) as they drop metrics to near-uniform values for both modalities, making it difficult to compare.
@@
Thermal 0.018 0.007 0.022 0.204
Color (grayscale) 0.538 0.536 0.531 0.601
Color (col. jitter) 0.800 0.792 0.801 0.841
refers to AP@0.5::0.95, following MSCOCO [40]
@@
Method Color Color (gray) Color (jitter) Thermal
SAM [33] + GT ann. 0.704 0.698 0.700 0.653
Grounded SAM [43] 0.380 0.351 0.361 0.193
OV-Seg [38] 0.362 0.340 0.365 0.217
ODISE [77] 0.504 0.450 0.486 0.232
Most frequent ground truth class label in a SAM mask assigned as label for entire mask
We quantify Segment Anything’s (SAM) [33] zero-shot segmentation sensitivity to modality change by measuring the spatial alignment of its mask outputs. We apply it to the RGB set and the 8-bit thermal image set (Sec. 3.2.2) with the default 32$\times$32 grid points to segment everything (Fig. 7). We compute the average precision (AP) of the predicted thermal masks, using the predicted RGB masks as ground truth, and compare the thermal AP to the APs of a color-jittered RGB set and a grayscale set. We find that SAM is sensitive to photometric changes, with large performance drops from color-jitter to grayscale to thermal (Tab. 7(a)).
@@
Finally, we create semantic SAM masks by assigning class labels to SAM instances based on the most frequent ground truth class within an instance (Fig. 7). We find that these masks attain much higher mIoU score (0.653) compared to the other zero-shot methods. Based on the results (Tab. 7(b)), we note that SAM can find object boundaries in the thermal domain reasonably well but can oversegment, resulting in poor AP scores when compared to RGB (which can also oversegment). When taken with ground truth semantics, this behavior is hidden.
@@ 4.5.2 Transfer learning with urban thermal datasets
None ImageNet Therm. Urban
@@
EfficientViT 0.687 0.725 0.714 4.8
ResNet18 0.702 0.713 0.706 12.3
ResNet50 0.682 0.713 0.728 26.7
ConvNext-B 0.625 0.717 0.697 89.4
@@
We test if pretraining on thermal urban data can assist transfer to field settings. We sample 40k unlabeled thermal images from [17, 71, 67, 11, 41, 22, 61, 76] and pretrain various networks using MoCoV2 [9]. We then finetune the pretrained networks on our training set and evaluate on our test set. Overall, we find sparse evidence suggesting any advantage of pretraining with existing urban thermal datasets (Tab. 8). Off-the-shelf ImageNet weights can provide strong performance in the thermal field domain, especially over training from scratch, without the effort and costs of pretraining.
@@ 5 Conclusion
We presented the Caltech Aerial RGB-Thermal Dataset, the first dataset tailored towards advancing thermal semantic perception and motion tracking algorithms in natural environments. We created four benchmarks encompassing semantic segmentation, image translation, and motion tracking, and demonstrate challenges faced by current thermal perception and localization algorithms. Semantic segmentation and image translation methods were notably affected by geographical domain shifts, reflecting the wide intra-class and inter-scenery distributions in our dataset, and temporal domain shifts, due to thermal inversion and crossover. Motion tracking in thermal suffered from feature sparsity, a challenge commonly encountered in real-world robotic deployments [1]. Our motion tracking benchmark penalizes current VIO/SLAM algorithms for assuming ideal feature-tracking conditions, serving as a testbed to fuel improvements in this area. Our dataset and benchmarks can be used into the future to help develop algorithms that expand the operating domains of robotics and computer vision.
Acknowledgements
This work was funded by the Office of Naval Research. We thank K. Koetje, K. Brodie, T. Hesser, T. Almeida, T. Cook, N. Spore, and R. Beach.